the sum of two numbers is 80. Three smaller numbers is larger than the bigger number by 48. Determine the value
Answer by stanbon(75887)

the sum of two numbers is 80. Three times the smaller number is larger than the bigger number by 48. Determine the value
----
Equations:
x + y = 80
3x - y = 48
---
Add to get:
4x = 128
x = 32 (smaller #)
---
solve for "y":
x + y = 80
32 + y = 80
y = 48 (larger #)
-----------------
Cheers,
Stan H.